OpenStudy (anonymous):

How To Write 44 49/80 As A Decimal.?

OpenStudy (texaschic101):

There is two ways to do this....you can turn the number into an improper fraction 44 49/80 = 3569/80 = 44.61 or you can temporarily disregard the whole number 44 and work with 49/80 which when divided = .61, then bring back the 44 and you have 44.61. Do you understand ?

OpenStudy (anonymous):

Kinda. I Never Learned Math When I Was Little So All This Is New To Me. But When I Typed It In, It Told Me I Was Wrong.

OpenStudy (texaschic101):

let me put it this way....you just divide. Look at this 3/10.....just divide 3 by 10 and you get your decimal .3. And like I said, if it is a mixed number, like your problem above, either turn it into an improper fraction and then divide or temporarily disregard the whole number, and divide the fraction, then add the whole number too it.

OpenStudy (texaschic101):

Try this one.......6/13.....what is the decimal ?

OpenStudy (anonymous):

0.42.?

OpenStudy (texaschic101):

no...try again.....6 divided by 13

OpenStudy (anonymous):

0.46

OpenStudy (texaschic101):

yes.....now try this mixed number.....2 1/3.....either turn it into an improper fraction or disregard the whole number and divide the fraction, then add it back to the whole number

OpenStudy (anonymous):

2.33

OpenStudy (texaschic101):

you got it...very good

OpenStudy (anonymous):

Thank You

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

\[44 \frac{49}{80}= \frac{44\times80+49}{80}= \frac{3520+49}{80}= \frac{3569}{80}=44.6125\] 2nd Method \[44 \frac{49}{80}= 44+ \frac{49}{80}= 44+0.6125=44.6125\]

OpenStudy (anonymous):

(:

OpenStudy (texaschic101):

dpasingh is exactly right :)